### v2

@@ Abstract
Due to their importance in various emerging applications, efficient algorithms for solving minimax problems have recently received increasing attention. However, many existing algorithms require prior knowledge of the problem parameters in order to achieve optimal iteration complexity. In this paper, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, to solve the smooth nonconvex-strongly concave and nonconvex-concave minimax problems respectively using a backtracking strategy, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$. Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ and $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$ respectively, where $\kappa$ is the condition number. As far as we know, the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm are the first completely parameter-free algorithms for solving nonconvex-strongly concave minimax problems and nonconvex-concave minimax problems respectively. Numerical results validate the efficiency of the proposed PF-AGP algorithm.
Keywords: minimax optimization problem; backtracking strategies; parameter-free alternating gradient projection algorithm; iteration complexity; machine learning
@@ 1 Introduction
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are nonempty closed and bounded convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function. Minimax problem is a critical class of optimization problems, known for its extensive research history and broad applications in areas such as machine learning, signal processing, and various other fields in recent years. Many practical problems can be formulated as in (P), including generative adversarial networks (GANs) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing (Lu et al., 2020), distributed nonconvex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben and Tal, 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh Abadeh et al., 2015) and several other applications.
@@
There are many research results on convex-concave minimax problems, i.e., $f(x,y)$ is convex with respect to $x$ and concave with respect to $y$. Nemirovski (Nemirovski, 2004) introduced the mirror-prox method, and Nestrov (Nesterov, 2007) proposed a dual extrapolation algorithm, both of which achieve an optimal iteration complexity of $\mathcal{O}(\varepsilon^{-1})$ for solving convex-concave minimax problems. For more details, see (Vandenberghe and Boyd, 2004; Chen et al., 2014, 2017; Dang and Lan, 2014; Monteiro and Svaiter, 2010; Nedić and Ozdaglar, 2009; Yang et al., 2020) and references therein.
Recently, various algorithms have been proposed for nonconvex-(strongly) concave minimax problems. For nonconvex-strongly concave minimax problems, several recent works (Jin et al., 2020; Lin et al., 2020a; Rafique et al., 2022; Lu et al., 2020) have proposed various algorithms that achieve a gradient complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{2}\varepsilon^{-2})$ for stationary points of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ (when $\mathcal{X}=\mathbb{R}^{n}$), or for stationary points of $f$, where $\kappa_{y}$ is the condition number for $f(x,\cdot)$. (Lin et al., 2020b) proposed an accelerated algorithm that improves the gradient complexity bound to $\tilde{\mathcal{O}}(\sqrt{\kappa_{y}}\varepsilon^{-2})$. Furthermore, Zhang et al. (Zhang et al., 2021) proposed a generic acceleration framework which can improve the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\varepsilon^{-2}\right)$.
@@
For general nonconvex-concave minimax problems, there are two types of algorithms, i.e., multi-loop algorithms and single-loop algorithms. Various multi-loop algorithms have been proposed in (Kong and Monteiro, 2021; Lin et al., 2020b; Nouiehed et al., 2019; Ostrovskii et al., 2021; Rafique et al., 2022; Thekumparampil et al., 2019). The best known iteration complexity of multi-loop algorithms for solving nonconvex-concave minimax problems is $\tilde{\mathcal{O}}(\varepsilon^{-2.5})$, which was achieved by (Lin et al., 2020b). For solving nonconvex-linear minimax problems, Pan et al. (Pan et al., 2021) proposed a new alternating gradient projection algorithm with the iteration complexity of $\mathcal{O}(\varepsilon^{-3})$. Various single-loop algorithms have also been proposed, e.g., the gradient descent-ascent (GDA) algorithm (Lin et al., 2020a), the hybrid block successive approximation (HiBSA) algorithm (Lu et al., 2020), the unified single-loop alternating gradient projection (AGP) algorithm (Xu et al., 2023), and the smoothed GDA algorithm (Zhang et al., 2020). Both the AGP algorithm and the smoothed GDA algorithm achieve the best known iteration complexity, i.e., $\mathcal{O}(\varepsilon^{-4})$, among single-loop algorithms for solving nonconvex-concave minimax problems.
@@
It should be noted that for most of the existing results mentioned above, achieving optimal complexity of the algorithm requires the assumption of knowing precise information about some parameters of the problem, such as the Lipschitz constant, or the strongly concave constant $\mu$, etc. Accurately estimating these parameters is often challenging, and conservative estimates can significantly impact algorithm performance (Lan et al., 2023). Therefore, designing parameter-free algorithms with complexity guarantees without relying on inputs of these parameters has attracted considerable attention recently.
@@
For nonconvex-strongly concave minimax problems, there are two main parameter-free algorithms: one based on AdaGrad stepsizes and the other based on backtracking. In the framework of AdaGrad stepsizes, Yang et al. (Yang et al., 2022) proposed NeAda, which can achieve convergence rates of $\tilde{\mathcal{O}}(\varepsilon^{-2})$ and $\tilde{\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively. Li et al. (Li et al., 2022) proposed TiAda, achieving convergence rates of ${\mathcal{O}}(\varepsilon^{-2})$ and ${\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively. In the framework of backtracking, Xu et al. (Xu et al., 2024) proposed a stochastic gradient descent ascent (GDA) method with backtracking (SGDA-B), achieving an upper bound on the total number of gradient calls of $\mathcal{O}(L\kappa^{2}\log(\kappa)\varepsilon^{-2})$ with $\mathcal{O}(L\kappa^{2}\varepsilon^{-2})$ gradient calls required per backtracking step. Meanwhile, Zhang et al. (Zhang et al., 2024) proposed a proximal alternating gradient descent ascent method (AGDA+), achieving an upper bound on the total number of gradient calls of $\mathcal{O}(L\kappa^{4}\varepsilon^{-2})$ with $\mathcal{O}(1)$ gradient calls required per backtracking step. Both methods do not require knowledge of the specific value of Lipschtiz constant $L$, but still require information about the strongly concave constant $\mu$ in the function. In the framework of backtracking, Xu et al. (Xu et al., 2024) proposed a stochastic GDA method with backtracking (SGDA-B) for nonconvex-concave minimax problems, achieving an upper bound on the total number of gradient calls of $\tilde{\mathcal{O}}(L^{3}\varepsilon^{-4})$ with $\mathcal{O}(L^{3}\varepsilon^{-4})$ gradient calls required per backtracking step.
@@ 1.1 Contributions
In this paper, based on the framework of backtracking, we propose two completely parameter-free alternating gradient projection algorithms, i.e., the PF-AGP-NSC algorithm and the PF-AGP-NC algorithm, for solving nonconvex-(strongly) concave minimax problems respectively, which does not require prior knowledge of parameters such as the Lipschtiz constant $L$ or the strongly concave constant $\mu$.
Moreover, we show that the total number of gradient calls of the PF-AGP-NSC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-strongly concave minimax problems is upper bounded by $\mathcal{O}\left(L\kappa^{3}\varepsilon^{-2}\right)$ where $\kappa$ is the condition number, while the total number of gradient calls of the PF-AGP-NC algorithm to obtain an $\varepsilon$-stationary point for nonconvex-concave minimax problems is upper bounded by $\mathcal{O}\left(L^{4}\varepsilon^{-4}\right)$. As far as we know, the PF-AGP-NSC algorithm is the first completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, and the PF-AGP-NC algorithm is the first completely parameter-free algorithm which achieves the best iteration complexity in single loop algorithms for solving nonconvex-concave minimax problems.
@@ 1.2 Organization
In Sect. 2, we propose two parameter-free alternating gradient projection (PF-AGP) algorithms for nonconvex-strongly concave and nonconvex-concave minimax problems, and we then analyze their corresponding gradient complexity under two different settings. We report some numerical results in Sect. 3 and make some conclusions in the last section.
@@
If $-f$ satisfies (1), $f(\cdot)$ is called $\theta$-strongly concave. Denote the sign function as $\mbox{sgn}(x)$, i.e.,
@@
2 Two completely parameter-free alternating gradient projection algorithms
In this section, we propose two completely parameter-free alternating gradient projection (PF-AGP) algorithms for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems (P) respectively. Both algorithms do not require prior knowledge of Lipschitz constants, and the algorithm under nonconvex-strong concave settings also does not require prior knowledge of strong concave constants. The two proposed algorithms are based on the AGP algorithm (Xu et al., 2023), which uses the gradient of a regularized version of the original function, i.e.,
$f_{k}(x,y)=f(x,y)-\frac{c_{k}}{2}\|y_{k}\|^{2},$
where $c_{k}\geq 0$ is a regularization parameters. At the $k$-th iteration, AGP algorithm consists of the following two gradient projection steps for updating both $x$ and $y$:
@@ 2 Two completely parameter-free alternating gradient projection algorithms
$y_{k+1}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\tfrac{1}{\gamma_{k}}\nabla_{y}{f}(x_{k},y_{k})-\frac{1}{\gamma_{k}}c_{k}y_{k}\right),$
@@
Although the AGP algorithm achieves the optimal iteration complexity among single loop algorithms for nonconvex-(strongly) concave minimax problems, it requires the knowledge of the Lipschtiz constant $L$ and the strongly concave coefficient $\mu$, or at least knowledge of the Lipschtiz constant $L$ to calculate the stepsizes $\beta_{k}$ and $\gamma_{k}$ under the nonconvex-strongly concave setting and the nonconvex-concave setting, respectively, which limits its applicability. To eliminate the need for the prior knowledge of the Lipschtiz constant $L$ or the strongly concave coefficient $\mu$, we propose two parameter-free AGP algorithms by employing backtracking strategies for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems, respectively. Similar backtracking strategies have been used in (Lan et al., 2023; Liu and Luo, 2022; Nesterov and Polyak, 2006; Nesterov, 2015).

### v3

@@ Abstract
Due to their importance in various emerging applications, efficient algorithms for solving minimax problems have recently received increasing attention. However, many existing algorithms require prior knowledge of the problem parameters in order to achieve optimal iteration complexity. In this paper, three completely parameter-free single-loop algorithms, namely PF-AGP-NSC algorithm, PF-AGP-NC algorithm and PF-AGP-NL algorithm, are proposed to solve the smooth nonconvex-strongly concave, nonconvex-concave minimax problems and nonconvex-linear minimax problems respectively using line search without requiring any prior knowledge about parameters such as the Lipschtiz constant $L$ or the strongly concave modulus $\mu$. Furthermore, we prove that the total number of gradient calls required to obtain an $\varepsilon$-stationary point for the PF-AGP-NSC algorithm, the PF-AGP-NC algorithm, and the PF-AGP-NL algorithm are upper bounded by  $\mathcal{O}\left(L^{2}\kappa^{3}\varepsilon^{-2}\right)$, $\mathcal{O}\left(\log^{2}(L)L^{4}\varepsilon^{-4}\right)$, and $\mathcal{O}\left(L^{3}\varepsilon^{-3}\right)$, respectively, where $\kappa$ is the condition number. To the best of our knowledge, PF-AGP-NC and PF-AGP-NL are the first completely parameter-free algorithms for solving nonconvex-concave and nonconvex-linear minimax problems, respectively. PF-AGP-NSC is a completely parameter-free algorithm for solving nonconvex-strongly concave minimax problems, achieving the best known complexity with respect to $\varepsilon$. Numerical results demonstrate the efficiency of the three proposed algorithms.
Keywords: minimax optimization; parameter-free alternating gradient projection algorithm; iteration complexity; machine learning
@@ 1 Introduction
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ and $\mathcal{Y}\subseteq\mathbb{R}^{m}$ are non-empty closed compact convex sets, and $f:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ is a smooth function. In recent years, the minimax problems have attracted the attention of scholars in the cross-disciplinary fields of optimization and machine learning due to its wide application. Many practical problems can be modeled as (P), such as generative adversarial networks (GAN) (Arjovsky et al., 2017; Goodfellow et al., 2014; Sanjabi et al., 2018), reinforcement learning (Dai et al., 2018), power control and transceiver design in signal processing Lu et al. (2020), distributed non-convex optimization (Giannakis et al., 2017; Giordano et al., 2018; Mateos et al., 2010), robust optimization (Ben-Tal et al., 2009; Gao and Kleywegt, 2023), statistical learning (Giordano et al., 2018; Shafieezadeh-Abadeh et al., 2015), etc. When $f(x,y)$ is (non)convex with respect to $x$ and (non)concave with respect to $y$, we call (P) the (non)convex-(non)concave minimax problem respectively.
@@
There are many research results on convex-concave minimax problems. The mirror neighbor method was proposed in (Nemirovski, 2004), and the dual extrapolation algorithm was proposed in (Nesterov, 2007), both of which achieve the optimal iteration complexity $\mathcal{O}(\varepsilon^{-1})$ for solving convex-concave minimax problems. For more related research results, please refer to (Vandenberghe and Boyd, 2004; Chen et al., 2014, 2017; Dang and Lan, 2014; Monteiro and Svaiter, 2010; Nedić and Ozdaglar, 2009; Yang et al., 2020) and their references.
@@
Recently, various algorithms have been proposed to solve the nonconvex-(strongly) concave minimax problems. For the nonconvex-strongly concave minimax problems, some recent studies (Jin et al., 2020; Lin et al., 2020a; Rafique et al., 2022; Lu et al., 2020) proposed various algorithms that achieve gradient complexity of $\tilde{\mathcal{O}}(\kappa_{y}^{2}\varepsilon^{-2})$ for stationary points of $\Phi(\cdot)=\max_{y\in\mathcal{Y}}f(\cdot,y)$ when $\mathcal{X}=\mathbb{R}^{n}$ or for stationary points of $f$ where $\kappa_{y}$ is the condition number of $f(x,\cdot)$. An acceleration algorithm was proposed in (Lin et al., 2020b), which improves the gradient complexity to $\tilde{\mathcal{O}}(\sqrt{\kappa_{y}}\varepsilon^{-2})$. In addition, Zhang et al. (Zhang et al., 2021) proposed a general acceleration framework, which improves the iteration complexity to $\mathcal{O}\left(\sqrt{\kappa_{y}}\varepsilon^{-2}\right)$. For general nonconvex-concave minimax problems, there are two types of algorithms, i.e., multi-loop algorithms and single-loop algorithms. Various multi-loop algorithms have been proposed in (Kong and Monteiro, 2021; Lin et al., 2020b; Nouiehed et al., 2019b; Ostrovskii et al., 2021; Rafique et al., 2022; Thekumparampil et al., 2019). The best known iteration complexity of multi-loop algorithms for solving nonconvex-concave minimax problems is $\tilde{\mathcal{O}}(\varepsilon^{-2.5})$, which was achieved by (Lin et al., 2020b). For solving nonconvex-linear minimax problems, Pan et al. (Pan et al., 2021) proposed a new alternating gradient projection algorithm with the iteration complexity of $\mathcal{O}(\varepsilon^{-3})$. Various single-loop algorithms have also been proposed for solving nonconvex-concave minimax problems, e.g., the gradient descent-ascent (GDA) algorithm (Lin et al., 2020a), the hybrid block successive approximation (HiBSA) algorithm (Lu et al., 2020), the unified single-loop alternating gradient projection (AGP) algorithm (Xu et al., 2023), and the smoothed GDA algorithm (Zhang et al., 2020). Both the AGP algorithm and the smoothed GDA algorithm achieve the best known iteration complexity, i.e., $\mathcal{O}(\varepsilon^{-4})$, among single-loop algorithms.
@@
It should be noted that for most of the above existing results, achieving the optimal complexity of the algorithm requires assuming accurate information about certain parameters of the known problem, such as the Lipschitz constant or the strong concavity modulus $\mu$, etc. Accurately estimating these parameters is usually challenging, and conservative estimates will significantly affect the algorithm performance (Lan et al., 2023). Therefore, designing parameter-free algorithms that do not rely on these parameter inputs and have complexity guarantees has attracted widespread attention in recent years.
@@
For nonconvex-strongly concave minimax problems, there are two main types of parameter-free algorithms: one based on AdaGrad steps and the other based on line search. In the framework of AdaGrad steps, the NeAda algorithm was proposed in (Yang et al., 2022), which obtains an $\varepsilon$-stationary point with an iteration complexity of $\tilde{\mathcal{O}}(\varepsilon^{-2})$ and $\tilde{\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively. The TiAda algorithm was proposed in (Li et al., 2022), which obtains an $\epsilon$-stationary point with an iteration complexity of ${\mathcal{O}}(\varepsilon^{-2})$ and ${\mathcal{O}}(\varepsilon^{-4})$ in deterministic and stochastic settings, respectively. Based on the line search framework, a stochastic gradient descent with backtracking (SGDA-B) algorithm was proposed in (Xu et al., 2024), which has an upper bound of $\mathcal{O}(L\kappa^{2}\log(\kappa)\varepsilon^{-2})$ on the total number of gradient calls to obtain an $\varepsilon$-stationary point. A proximal alternating gradient descent (AGDA+) method was proposed in (Zhang et al., 2025) which achieves an upper bound of the total number of gradient calls of $\mathcal{O}(\bar{L}^{2}\bar{\kappa}^{4}\log(\mathcal{R})\varepsilon^{-2})$, where $\bar{\kappa}=\bar{L}/\bar{\mu}$, $\bar{L}=\mathcal{R}\tilde{l}$ and $\bar{\mu}=\tilde{\mu}/\mathcal{R}$, $\mathcal{R}=\max\{\tilde{\mu}/\mu,L/\tilde{l},1\}$, $\tilde{\mu}\geq\mu$ and $\tilde{l}\leq L$ denote the initial estimate of the concavity modulus $\mu$ and the local Lipschitz constant $L$.
For the nonconvex-concave minimax problems, a stochastic GDA method with backtracking (SGDA-B) is proposed in (Xu et al., 2024), which obtains an upper bound of $\tilde{\mathcal{O}}(L^{3}\varepsilon^{-4})$ on the total number of gradient calls to obtain an $\varepsilon$-stationary point.
For nonconvex-linear minimax problems (a special class of nonconvex-concave minimax problems), before the results of this paper, there were no relevant results on parameter-free optimization algorithms with an iteration complexity of $\mathcal{O}(\varepsilon^{-3})$.
@@ 1.1 Contributions
Based on the line search framework, this paper proposes three completely parameter-free alternating gradient projection algorithms, namely PF-AGP-NSC algorithm, PF-AGP-NC algorithm and PF-AGP-NL algorithm, which are used to solve nonconvex-strongly concave, nonconvex-concave and nonconvex-linear minimax problems respectively. These algorithms do not require prior knowledge of any problem-related parameters, including the Lipschtiz constant $L$ and the strongly concave modulus $\mu$. Furthermore, we prove that the total number of gradient calls required to obtain an $\varepsilon$-stationary point for the PF-AGP-NSC algorithm, the PF-AGP-NC algorithm, and the PF-AGP-NL algorithm are upper bounded by  $\mathcal{O}\left(L^{2}\kappa^{3}\varepsilon^{-2}\right)$, $\mathcal{O}\left(\log^{2}(L)L^{4}\varepsilon^{-4}\right)$, and $\mathcal{O}\left(L^{3}\varepsilon^{-3}\right)$, respectively.
Table 1 provides a detailed comparison of the three proposed parameter-free optimization algorithms with existing parameter-free optimization algorithms for solving nonconvex-strongly concave (NC-SC), nonconvex-concave (NC-C), and nonconvex-linear (NC-L) minimax problems. Column “Completely PF” indicates whether the algorithm is completely parameter-free (yes/no), and column “Prior Information” specifies any prior knowledge or parameters required by the algorithm, where $\underline{S}=\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}f(x,y)$ and $\mathcal{D}_{y}=\max\{\|y_{1}-y_{2}\|\mid y_{1},y_{2}\in\mathcal{Y}\}$ and ”None” indicates that no prior information is required. In “Gradient Complexity” column, $\bar{\kappa}\geq\kappa$, $\mathcal{R}$ is a constant which depends on $L$ and $\mu$ and $\mathcal{R}\leq\kappa$.
To the best of our knowledge, the PF-AGP-NL algorithm is the first completely parameter-free single-loop algorithm for solving nonconvex-linear minimax problems, achieving the best known gradient complexity within the single-loop algorithms. The PF-AGP-NSC and the PF-AGP-NC algorithm is the first completely parameter-free single-loop algorithm for solving nonconvex-strongly concave and nonconvex-concave minimax problems, achieving the best known gradient complexity with respect to $\varepsilon$ with the existing single-loop algorithms respectively.
It should be pointed out that the complexity of the proposed PF-AGP-NSC and PF-AGP-NC algorithms is slightly worse than that of the SGDA-B algorithm in terms of the Lipschitz constant L and the condition number $\kappa$, but the SGDA-B algorithm requires some prior information about the remaining parameters and is not a completely parameter-free algorithm. If some other prior information is also allowed, the complexity of the proposed algorithm can also be improved. To better illustrate this point, if two additional parameters $\mu$ and $\underline{S}$ are known in advance, in Appendix A we propose a restarted variant of the PF-AGP-NSC algorithm, named rPF-AGP-SC, and prove that the gradient complexity of the algorithm is improved to $\mathcal{O}\left(\log(L)L\kappa^{2}\varepsilon^{-2}\right)$, which is better than that of the SGDA-B.
Problems Algorithms Gradient Complexity Completely PF Prior Information
NC-SC NeAda $\mathcal{O}(\log(\kappa)\kappa^{4.5}L^{4}\varepsilon^{-2})$ Yes None
TiAda $\mathcal{O}(\kappa^{10}\varepsilon^{-2})$ Yes None
SGDA-B $\mathcal{O}(L\kappa^{2}\log(\kappa)\varepsilon^{-2})$ No $\mu$, $\underline{S}$
AGDA+ $\mathcal{O}(\bar{L}^{2}\bar{\kappa}^{4}\log(\mathcal{R})\varepsilon^{-2})$ Yes None
PF-AGP-NSC $\mathcal{O}(L^{2}\kappa^{3}\varepsilon^{-2})$ Yes None
NC-C SGDA-B $\mathcal{O}(L^{3}\log({L}{\varepsilon^{-1}})\varepsilon^{-4})$ No $\mathcal{D}_{y}$
PF-AGP-NC $\mathcal{O}(\log^{2}(L)L^{4}\varepsilon^{-4})$ Yes None
NC-L PF-AGP-NL $\mathcal{O}(L^{3}\varepsilon^{-3})$ Yes None
Table 1: Comparison of Gradient Complexity of Existing Algorithms for Nonconvex-Concave Minimax Problems.
@@ 1.2 Organization
In Section 2, we propose two parameter-free single-loop algorithms, PF-AGP-NSC and PF-AGP-NC, for the nonconvex-strongly concave and nonconvex-concave minimax problems, and analyze their corresponding gradient complexities in two different settings. In Section 3, we propose a parameter-free alternating gradient projection (PF-AGP-NL) algorithm for the nonconvex-linear minimax problem and analyze its corresponding gradient complexity. We report some numerical results in Section 4 and draw some conclusions in the last section.
@@
$f(\cdot)$ is called $\theta$-strongly concave if $-f$ satisfies (1). The symbolic function is represented as $\mbox{sgn}(x)$, that is:
@@
2  Completely parameter-free single-loop algorithms for nonconvex- (strongly) concave setting
In this section, we propose two completely parameter-free single-loop algorithms for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems (P) respectively. Both algorithms do not require prior knowledge of Lipschitz constants, and the algorithm under nonconvex-strong concave settings also does not require prior knowledge of strong concavity coefficient. The two proposed algorithms are based on the AGP algorithm (Xu et al., 2023), which uses the gradient of a regularized version of the original function, i.e.,
$f_{k}(x,y)=f(x,y)-\frac{c_{k}}{2}\|y\|^{2},$
where $c_{k}\geq 0$ is a regularization parameter. At the $k$-th iteration, AGP algorithm consists of the following two gradient projection steps for updating both $x$ and $y$:
@@ 2  Completely parameter-free single-loop algorithms for nonconvex- (strongly) concave setting
$y_{k+1}$ $=\mathcal{P}_{\mathcal{Y}}\left(y_{k}+\tfrac{1}{\gamma_{k}}\nabla_{y}{f}(x_{k+1},y_{k})-\frac{1}{\gamma_{k}}c_{k}y_{k}\right),$
@@
Although the AGP algorithm achieves the optimal iteration complexity dependence on $\varepsilon$ among single loop algorithms for nonconvex-(strongly) concave minimax problems, it requires the knowledge of the Lipschtiz constant $L$ and the strongly concave coefficient $\mu$, or at least knowledge of the Lipschtiz constant $L$ to calculate the stepsizes $\beta_{k}$ and $\gamma_{k}$ under the nonconvex-strongly concave setting and the nonconvex-concave setting, respectively, which limits its applicability. To eliminate the need for the prior knowledge of the Lipschtiz constant $L$ or the strongly concave coefficient $\mu$, we propose two parameter-free AGP algorithms by employing backtracking strategies for solving the nonconvex-strongly concave and the nonconvex-concave minimax problems, respectively. Similar backtracking strategies have been used in (Lan et al., 2023; Liu and Luo, 2022; Nesterov and Polyak, 2006; Nesterov, 2015).